OpenStudy (anonymous):

a 90 in pipe is cut into two pieces. One piece is 20 in more than the other. Find the length of each piece?

OpenStudy (anonymous):

is it 65 and 25?

OpenStudy (anonymous):

representation: 1 piece= x 2 piece= x+20 equation: x+(x+20) = 90 x+x+20 = 90 2x = 90-20 x = 70/2 x = 35 therefor: the parts are x+20= 35+20 =55 x =35

OpenStudy (anonymous):

so its 55 and 35?

OpenStudy (anonymous):

yeah. .